# Computing sessile droplet shapes on arbitrary surfaces with a new pairwise force smoothed particle hydrodynamics model

### v1

@@ 1 Introduction
Since its inception, SPH has been applied to a wide variety of problems, including astrophysics (Springel, 2010), coastal engineering (Barreiro et al., 2013), porous media flow (Zhu and Fox, 2001), and even blood flow in injured arteries (Müller et al., 2004). The key advantage of SPH in simulating droplet behaviour is that the scheme does not necessitate the tracking of the sharp interface between the liquid and the air. Indeed, it does not track the interface directly at all; instead, the interface is deduced a posteriori based on the density field. This approach provides the needed flexibility to model droplets with non-trivial shapes on complex substrates that would otherwise pose a significant challenge for interface tracking or fixed grid methods (Ye et al., 2019). In the spirit of this interface-free approach, in the past decade a modified SPH method has emerged in which the Young-Laplace pressure boundary condition at the free surface is replaced with inter-particle forces that mimic cohesion and adhesion, in what is known as the pairwise force SPH method (PF-SPH).
@@
Kordilla et al. (2013), Tartakovsky and Panchenko (2016), and Shigorina et al. (2017) all used a PF-SPH method to study droplet shapes on rough surfaces, although each used a different function for the pairwise force. There is, in general, a lack of consensus regarding the best pairwise force formulation, and a lack of understanding about what effect the formulation has on the simulated surface tension and contact angle. Existing pairwise force profiles in the literature are rarely physically motivated and seldom investigated in their own right. With this in mind, our main contribution is to propose a new force profile for PF-SPH that is physically motivated, scale-independent, and carefully validated through the reproduction of multiple independent physical phenomena. To this end, we have developed a new method for calibrating the contact angle on an ideal surface by fitting whole droplet shapes obtained from semi-analytical solutions to the Young-Laplace equation. With the new PF-SPH model thoroughly calibrated and validated, we then demonstrate its application to several important test problems from the literature. First, we calculate the equilibrium shape of settling droplets on microscopically rough and chemically patterned substrates. Then, we apply the model to simulate a scenario from an agricultural application: a droplet impacting and settling on a virtual plant leaf.
@@ 2.2 Discretisation
$w(\tau)$ =$=\begin{cases}(1-\tau)^{4}(4\tau+1),&0\leq\tau<1,\\ 0,&\tau\geq 1.\end{cases}$
@@ 2.2.4 Solid boundary treatment
At the fluid-solid interface, we have the no-slip condition: $\bm{\mathrm{v}}=\bm{\mathrm{0}}$. We impose this condition indirectly, representing the solid substrate with fixed dummy particles, initialised on a regular grid with spacing $\Delta x$ (Macia et al., 2011). Multiple layers of these particles are used to avoid an SPH neighbourhood deficiency at the boundary. The solid particles have the same physical properties as the fluid at rest, with mass $\rho_{0}\Delta x^{3}$ and density $\rho_{0}$, but with zero velocity. The solid particles are included in the summations over $j$ in the discrete continuity and momentum equations (3) and (4), as if they were fluid particles. Pressure forces and the repulsion due to pairwise forces ensure that fluid particles do not penetrate the boundary. Figure 1 shows a typical setup in which a droplet of fluid particles is about to impact a flat solid boundary.
@@ 2.2.5 Time integration
Calculate $\text{Calculate }\bm{\mathrm{A}}^{\left(k\right)}\text{ and }Q^{\left(k\right)},$
@@
Calculate $\text{Calculate }\bm{\mathrm{A}}^{\left(k+\frac{1}{2}\right)}\text{ and }Q^{\left(k+\frac{1}{2}\right)},$
@@ 2.3 Implementation details
for each fluid particle $i$. Any sum over $j=1,\dots,N$ then becomes a sum over $j\in\mathcal{J}(i)$. Since density fluctuations are small in this weakly compressible scheme, the number of neighbours is approximately constant (around $34\pi\kappa^{3}/$ 3). Thus, the complexity of calculating the particle interactions becomes $\mathcal{O}(N)$.
@@ 2.4 Pairwise force model for interfacial phenomena
$f_{\mathrm{ff}}(1.1\Delta x/H)=0,\quad f_{\mathrm{fs}}(2\Delta x/H)=0.$ (repulsion-to-attraction point)
The last constraint, controlling the zero crossing of the force profile, is intentionally different for the fluid-fluid ($\mathrm{ff}$) and the fluid-solid ($\mathrm{fs}$) interactions. The fluid-fluid force profile is attractive at a shorter distance than the fluid-solid to prioritise cohesion over adhesion for stability at the contact line. Fourth-degree polynomials are simple to construct from these five constraints, and so we have:
$f_{ij}(\tau)=\begin{cases}0,&\tau>1,\\ 1-23.5\tau^{2}+43\tau^{3}-20.5\tau^{4},&\mathrm{labels}(i,j)=\mathrm{(fluid,fluid)},\\ 1-11\tau^{2}+18\tau^{3}-8\tau^{4},&\mathrm{labels}(i,j)=\mathrm{(fluid,solid)}.\end{cases}$ (8)
Figure 2 shows these force profiles over the range of possible particle distances.
Figure 2: Distance-dependent pairwise force profiles $f_{\mathrm{ff}}$ and $f_{\mathrm{fs}}$ over the kernel support, $\tau\in[0,1]$ or $\|\bm{\mathrm{x}}_{ij}\|/\Delta x\in[0,\kappa]$. Positive values indicate repulsion and negative values indicate attraction. Note that we have plotted the force profiles with respect to the particle width $\Delta x$ rather than the dimensionless $\tau$, to highlight the physically motivated choice of zero crossing discussed in Section 4.4.
@@ 3.1 Laplace pressure
where $d$ is the spatial dimension (taken here to be $d=3$), $V(r)$ is the volume of a sphere of radius $r$, and $\bm{\mathrm{F}}_{ij}$ is the sum of the pressure force and pairwise force that particle $i$ experiences due to particle $j$. The outer summation (over $i$) includes only particles in the set $\mathcal{I}(r)$ of particles within a distance $r$ of the centre of the droplet. This so-called virial radius $r<R$ is used in place of the actual droplet radius $R$ to exclude the region near the interface (Tartakovsky and Meakin, 2005). When $r=R$, we see a divergence of the pressure near the free surface due to the neighbourhood deficiency in the SPH approximations there. Figure 3 shows the virial pressure measured at different radii, with the pressure clearly diverging as $r$ approaches $R$. For accurate estimates of the virial pressure we take an average of $P(r)$ over the interval $r\in[R-3H,R-2H]$. If $R\leq 4H$, we take $r=R-2H$.
@@
Radius, $R$ $[0.8,1.2]\,\mathrm{mm}$
$s_{\mathrm{ff}}$ $[0,2\cdot 10^{-3}]\,\mathrm{N/m}$
@@ 3.2 Oscillating droplets
With material properties given in Table 2, we initialise a spherical droplet of radius $0.788$mm with particles that we randomly perturb by no more than $0.1\Delta x$, and allow the particle distribution to settle for 1ms.
@@ 4.1 Microstructured surface
Figure 13 shows a comparison between two almost identical numerical experiments – the only difference being the initial kinetic energy of the droplet before ‘impact’ with the surface. Figure 13(a) shows a settled droplet that was initialised with zero velocity. This droplet sits atop the pillars on the substrate, without enough energy to infiltrate the gaps between them. The droplet in Figure 13(b) was initialised with a velocity in the vertical direction of $-0.1\,\mathrm{m/s}$, allowing it to overcome the energy barrier discussed by Dupuis and Yeomans (2005) and transition from a ‘suspended’ to a ‘collapsed’ state, in which the fluid has infiltrated between the pillars. That the present model reproduces this qualitative behaviour suggests it should be applicable to problems with complex surfaces.
@@
# Particles 113 000113000113\,000113 000 fluid particles
@@ 4.2 Chemically patterned surface
[…]
@@
# Particles 164 968164968164\,968164 968 fluid particles
@@ 4.3 Wheat leaf
[…]
@@
# Particles 267 731267731267\,731267 731 fluid particles
@@ 4.4 Discussion on pairwise force profiles
The exact design of the pairwise force profile $f_{ij}$ is in need of more investigation before the PF-SPH model can be applied as readily as, for example, the continuum surface force formulation. Literature on the effect of the force profile is scarce. Tartakovsky and Panchenko (2016) report some analytical results predicting the surface tension from $f_{ij}(\tau)$ and $s_{ij}$ in multiphase simulations, but we could not verify these results in our single phase simulations.
We have designed the fluid-fluid pairwise force curve to have its zero crossing at approximately the ‘rest distance’ of the particles. To do this, we imagined the particles as spheres, packed optimally in a face-centred cubic layout. The Voronoi diagram of this arrangement is the rhombic dodecahedral honeycomb. If we equate the volume of one of our particles of diameter $\Delta x$ with the volume of a rhombic dodecahedron, we find that the distance between neighbouring particles in such a packing is approximately $1.1\Delta x$. This was the geometric motivation behind the location of the zero crossing of the force profile in (7).
Clearly, however, this will lead to most fluid-fluid pairwise force interactions being attractive, which could cause excess numerical stress in the fluid. In the simulations tested, the mostly-attractive pairwise force has the somewhat desirable effect of keeping the particle distribution ordered – for example, in a high velocity impact event – which ensures that interpolation error is low. More investigation is needed to quantify the dissipative effect of the pairwise forces before this model is applied to viscosity-dependent scenarios.
@@ 5 Conclusion
The test cases we present in Section 4 demonstrate that this method should be applicable to a broad range of droplet-related problems on substrates of interest, at least for sessile droplets and low-inertial flows. In future work, we intend to model the chemical heterogeneity of plant leaf surfaces by varying the adhesive parameter $s_{\mathrm{fs}}$ across the surface to more accurately reflect the real surface chemistry. Other potential applications include the study of spreading and impaction on rough or patterned surfaces. With careful parameter calibration, the pairwise force model shows promise for simulating droplets on otherwise challenging substrates.
@@ Acknowledgements
This work was financially supported by an ARC Research Training Program scholarship, as well as the ARC Linkage Grant LP160100707 and associated industry partners Syngenta and Nufarm.

### v2

@@ 1 Introduction
Since its inception, SPH has been applied to a wide variety of problems, including astrophysics (Springel, 2010), coastal engineering (Barreiro et al., 2013), porous media flow (Zhu and Fox, 2001), and even blood flow in injured arteries (Müller et al., 2004). The key advantage of SPH in simulating droplet behaviour is that the scheme does not necessitate the tracking of the sharp interface between the liquid and the air. Indeed, it does not track the interface directly at all; instead, the interface is deduced a posteriori based on the density field.
Some practitioners of SPH introduce surface tension effects by means of the continuum surface force (CSF). This idea dates back to Brackbill et al. (1992) (in general) and Morris (2000) (applied to SPH). The CSF method starts with the Young-Laplace pressure boundary condition at the liquid-gas interface, but translates the normal force per unit area on the interface into a force per unit volume throughout the liquid, by means of a smoothing function. That is, a force $\bm{\mathrm{F}}_{\mathrm{CSF}}$ is introduced everywhere in the fluid:
$\bm{\mathrm{F}}_{\mathrm{CSF}}=\delta_{\mathrm{CSF}}(\bm{\mathrm{x}})\,\sigma K(\bm{\mathrm{x}})\hat{\bm{\mathrm{n}}}(\bm{\mathrm{x}}),$
where $\delta$ is the smoothing function (maximised at the interface, and vanishing away from it), $\sigma$ is the surface tension, $K=-\nabla\cdot\hat{\bm{\mathrm{n}}}$ is the curvature of the interface, and $\hat{\bm{\mathrm{n}}}(\bm{\mathrm{x}})$ is the unit normal of the interface, extended into the fluid (Morris, 2000). The success of these methods hinges on robust calculation of the smoothing function and the unit normal ‘field’. Recent approaches (Vergnaud et al., 2022; Blank et al., 2024; Antuono et al., 2025) require kernel corrections, which increase the computational burden of the method significantly, and smoothing of the normal field, as well as special corrections for thin jets. The matter is further complicated when one considers contact angle effects for droplets on substrates – Chiron et al. (2019) resort to tracking the liquid-solid interface explicitly. This additional complexity yields accurate results; however, our aim is to produce a robust method applicable to droplet simulations on arbitrarily rough substrates. As such, we seek a method that is more computationally efficient and robust to interfaces with high curvatures, which we expect to be prevalent in our application.
We choose the SPH method because it provides the needed flexibility to model droplets with non-trivial shapes on complex substrates that would otherwise pose a significant challenge for interface tracking or fixed grid methods (Ye et al., 2019). In the spirit of this interface-free approach, in the past decade a modified SPH method has emerged in which the Young-Laplace pressure boundary condition at the free surface is replaced with inter-particle forces that mimic cohesion and adhesion, in what is known as the pairwise force SPH method (PF-SPH).
@@
Kordilla et al. (2013), Tartakovsky and Panchenko (2016), and Shigorina et al. (2017) all used a PF-SPH method to study droplet shapes on rough surfaces, although each used a different function for the pairwise force. Santacruz-Yunga et al. (2025) use a potential based on the popular Lennard-Jones potential for the pairwise interactions, which in this framework is analogous to a pairwise force. There is, in general, a lack of consensus regarding the best pairwise force formulation, and a lack of understanding about what effect the formulation has on the simulated surface tension and contact angle. Existing pairwise force profiles in the literature are rarely physically motivated and seldom investigated in their own right. Nevertheless, the PF-SPH method is very attractive for its simplicity, computational efficiency, and robustness to thin features of both substrates and droplets. With this in mind, our main contribution is to propose a new force profile for PF-SPH that is physically motivated, scale-independent, and carefully validated through the reproduction of multiple independent physical phenomena. To this end, we have developed a new method for calibrating the contact angle on an ideal surface by fitting whole droplet shapes obtained from semi-analytical solutions to the Young-Laplace equation. With the new PF-SPH model thoroughly calibrated and validated, we then demonstrate its application to several important test problems from the literature. First, we calculate the equilibrium shape of settling droplets on microscopically rough and chemically patterned substrates. Then, we apply the model to simulate a scenario from an agricultural application: a droplet impacting and settling on a virtual plant leaf.
@@ 2.2 Discretisation
$w(\tau)$ $=\begin{cases}(1-\tau)^{4}(4\tau+1),&0\leq\tau<1,\\ 0,&\tau\geq 1.\end{cases}$
@@ 2.2.4 Solid boundary treatment
At the fluid-solid interface, we have the no-slip condition: $\bm{\mathrm{v}}=\bm{\mathrm{0}}$. We impose this condition indirectly, representing the solid substrate with fixed dummy particles, initialised on a regular grid with spacing $\Delta x$. Multiple layers of these particles are used to avoid an SPH neighbourhood deficiency at the boundary. The solid particles have the same physical properties as the fluid at rest, with mass $\rho_{0}\Delta x^{3}$ and density $\rho_{0}$, but with zero velocity. This simple approximation of the no-slip condition has been shown to be consistent in the limit as $H\rightarrow 0$, although does not reproduce it exactly in simulations. Other schemes utilise mirroring approaches (Macia et al., 2011), in which fluid properties are interpolated and extended into the solid boundary using surface normals. In our study of droplets on rough surfaces, however, this approach would be quite complex to implement for a general surface. For example, it is not straightforward to map any point in space to its closest surface point and thus the appropriate surface normal. Furthermore, some boundary points may be assigned incorrect properties because their associated interpolation point above the solid is in the gas phase, which we do not discretise with SPH particles. The subject of more accurate boundary condition treatment is thus left to future work, and the interested reader is referred to Macia et al. (2011) for more details.
The “dummy” solid particles are included in the summations over $j$ in the discrete continuity and momentum equations (3) and (4), as if they were fluid particles. Pressure forces and the repulsion due to pairwise forces ensure that fluid particles do not penetrate the boundary. Figure 1 shows a typical setup in which a droplet of fluid particles is about to impact a flat solid boundary.
@@ 2.2.5 Time integration
$\text{Calculate }\bm{\mathrm{A}}^{\left(k\right)}\text{ and }Q^{\left(k\right)},$
@@
$\text{Calculate }\bm{\mathrm{A}}^{\left(k+\frac{1}{2}\right)}\text{ and }Q^{\left(k+\frac{1}{2}\right)},$
@@ 2.3 Implementation details
for each fluid particle $i$. Any sum over $j=1,\dots,N$ then becomes a sum over $j\in\mathcal{J}(i)$. Since density fluctuations are small in this weakly compressible scheme, the number of neighbours is approximately constant (around $4\pi\kappa^{3}/3$). Thus, the complexity of calculating the particle interactions becomes $\mathcal{O}(N)$.
@@ 2.4 Pairwise force model for interfacial phenomena
$f_{ij}(\kappa^{*}\Delta x/H)=0.$ (repulsion-to-attraction point)
The last constraint, controlling the zero crossing of the force profile, is intentionally different for the fluid-fluid ($\mathrm{ff}$) and the fluid-solid ($\mathrm{fs}$) interactions. We choose $\kappa^{*}$ to be $1.1$ and $2$ for the fluid-fluid and the fluid-solid interactions, respectively. The fluid-fluid force profile is attractive at a shorter distance than the fluid-solid to prioritise cohesion over adhesion for stability at the contact line.
We have chosen the zero crossing of the fluid-fluid pairwise force curve to be approximately the ‘rest distance’ of the particles. To do this, we imagine optimally packed spheres in a face-centred cubic layout. The Voronoi diagram of this arrangement is the rhombic dodecahedral honeycomb. If we equate the volume of one of our particles ($(\Delta x)^{3}$) with the volume of a rhombic dodecahedron, we find that the distance between neighbouring particles in such a packing is approximately $1.1\Delta x$. This is the geometric motivation behind the location of the zero crossing of the force profile in (7).
The minimum degree of a polynomial satisfying our five constraints is four, and so we have:
$f_{ij}(\tau)=\begin{cases}1+\alpha_{ij}\tau^{2}-2(2+\alpha_{ij})\tau^{3}+(3+\alpha_{ij})\tau^{4},&0\leq\tau\leq 1,\\ 0,&\tau>1,\end{cases}$ (8)
where we use $\alpha_{\mathrm{ff}}=-23.5$ for fluid-fluid interactions, and $\alpha_{\mathrm{fs}}=-11$ for fluid-solid interactions, to satisfy the zero crossing constraints above. Figure 2 shows these force profiles over the range of possible particle distances.
Figure 2: Distance-dependent pairwise force profiles $f_{\mathrm{ff}}$ and $f_{\mathrm{fs}}$ over the kernel support, $\tau\in[0,1]$ or $\|\bm{\mathrm{x}}_{ij}\|/\Delta x\in[0,\kappa]$. Positive values indicate repulsion and negative values indicate attraction. Note that we have plotted the force profiles with respect to the particle width $\Delta x$ rather than the dimensionless $\tau$, to highlight the physically motivated choice of zero crossing discussed in Section $4.$
@@
While the PF-SPH model may seem an oversimplification of the complex (and small-scale) behaviour that is molecular cohesion and adhesion, it is best thought of as a coarse approximation of these phenomena, at the resolution of the smoothed particle hydrodynamics discretisation. We will show in the next section that the resulting surface tension is linearly related to the pairwise force strength $s_{\mathrm{ff}}$, and that this simple model captures complex behaviour in a predictable way.
@@ 3.1 Laplace pressure
where $d$ is the spatial dimension (taken here to be $d=3$), $V(r)$ is the volume of a sphere of radius $r$, and $\bm{\mathrm{F}}_{ij}$ is the sum of the pressure force and pairwise force that particle $i$ experiences due to particle $j$. The outer summation (over $i$) includes only particles in the set $\mathcal{I}(r)$ of particles within a distance $r$ of the centre of the droplet. This so-called virial radius $r<R$ is used in place of the actual droplet radius $R$ to exclude the region near the interface (Tartakovsky and Meakin, 2005). When $r=R$, we see a divergence of the pressure near the free surface due to the neighbourhood deficiency in the SPH approximations there. Figure 3 shows the virial pressure measured at different radii, with the pressure clearly diverging as $r$ approaches $R$. For accurate estimates of the virial pressure we take an average of $P(r)$ over the interval $r\in[R-3H,R-2H]$. If $R\leq 4H$, we take r=R2Hr=R-2H.
@@
Radius, $R$ $0.8,1.2]\,\mathrm{mm}$
$s_{\mathrm{ff}}$ $0,2\cdot 10^{-3}]\,\mathrm{N/m}$
@@ 3.2 Oscillating droplets
With material properties given in Table 2, we initialise a spherical droplet of radius 0.7880.788mm with particles that we randomly perturb by no more than $0.1\Delta x$, and allow the particle distribution to settle for 1ms.
@@ 4.1 Microstructured surface
Figure 13 shows a comparison between two almost identical numerical experiments – the only difference being the initial kinetic energy of the droplet before ‘impact’ with the surface. Figure 13(a) shows a settled droplet that was initialised with zero velocity. This droplet sits atop the pillars on the substrate, without enough energy to infiltrate the gaps between them. The droplet in Figure 13(b) was initialised with a velocity in the vertical direction of $0.1\,\mathrm{m/s}$, allowing it to overcome the energy barrier discussed by Dupuis and Yeomans (2005) and transition from a ‘suspended’ to a ‘collapsed’ state, in which the fluid has infiltrated between the pillars. That the present model reproduces this qualitative behaviour suggests it should be applicable to problems with complex surfaces.
@@
# Particles 113 000113\,000 fluid particles
@@ 4.2 Chemically patterned surface
$s_{\mathrm{fs}}$ $7\text{ (hydrophilic)},\;2.8\text{ (hydrophobic)}\,\mathrm{mN/m}$
@@
# Particles 164 968164\,968 fluid particles
@@ 4.3 Wheat leaf
$s_{\mathrm{fs}}$ $5.6\text{ (leaf)},\;7.875\text{ (hairs)}\,\mathrm{mN/m}$
@@
# Particles 267 731267\,731 fluid particles
@@ 4.4 Discussion on pairwise force profiles
The exact design of the pairwise force profile $f_{ij}$ is in need of more investigation before the PF-SPH model can be applied as readily as, for example, the continuum surface force formulation. Literature on the effect of the force profile is scarce. Tartakovsky and Panchenko (2016) report some analytical results predicting the surface tension from $f_{ij}(\tau)$ and $s_{ij}$ in multiphase simulations, but we could not verify these results in our single phase simulations. Our force profile causes most fluid-fluid particle interactions to be attractive, which could cause excess numerical stress in the fluid. In the simulations tested, the mostly-attractive pairwise force has the somewhat desirable effect of keeping the particle distribution ordered – for example, in a high velocity impact event – which ensures that interpolation error is low. More investigation is needed to quantify the dissipative effect of the pairwise forces before this model is applied to viscosity-dependent scenarios.
@@ 5 Conclusion
The test cases we present in Section 4 demonstrate that this method should be applicable to a broad range of droplet-related problems on substrates of interest, at least for sessile droplets and low-inertial flows. In future work, we intend to model the chemical heterogeneity of plant leaf surfaces by varying the adhesive parameter $s_{\mathrm{fs}}$ across the surface to more accurately reflect the real surface chemistry. Other potential applications include the study of spreading and impaction on rough or patterned surfaces. With the careful parameter calibration developed here, the pairwise force model shows promise for simulating droplets on otherwise challenging substrates.
@@ Acknowledgements
This work was financially supported by an Australian Research Council Research Training Program scholarship, as well as the Australian Research Council Linkage Grant LP160100707 and associated industry partners Syngenta and Nufarm.